HETAV PATEL

Grade 11,

N men and n women sit along a line alternatively in x ways and along a circle in y ways such that x =10y,then the number of ways in which n men can sit at a round table so that all shall not have the same neighbours is, a)6 b)12 c)36 d)none

N men and n women sit along a line alternatively in x ways and along a circle in y ways such that x =10y,then the number of ways in which n men can sit at a round table so that all shall not have the same neighbours is,
a)6
b)12
c)36
d)none

Grade:12th pass

1 Answers

Anshumaan Paliwal
7 years ago
x/y = 2*n!*n! / [(n-1)!*n!] =2n
We have x=10y, so we get n=5.
Therefore, the required number is = 1/2 * 4! =12- b)